OpenStudy (anonymous):

aylor organized her rock collection intaylor organized her rock collection into boxes. She put the same number of rocks into each box. She made this chart to keep track of the total number of rocks she organized. What is the total number of rocks that Taylor has in 5 boxes? Taylor's Rocks Number of Boxes Total Number of Rocks 1 7 2 14 3 21 4 28 5 ? aylor organized her rock collection into boxes. She put the same number of rocks into each box. She made this chart to keep track of the total number of rocks she organized. What is the total number of rocks that Taylor has in 5 boxes? Taylor's Rocks Number of Boxes Total Number of Rocks 1 7 2 14 3 21 4 28 5 ? aylor organized her rock collection into boxes. She put the same number of rocks into each box. She made this chart to keep track of the total number of rocks she organized. What is the total number of rocks that Taylor has in 5 boxes? Taylor's Rocks Number of Boxes Total Number of Rocks 1 7 2 14 3 21 4 28 5 ? aylor organized her rock collection into boxes. She put the same number of rocks into each box. She made this chart to keep track of the total number of rocks she organized. What is the total number of rocks that Taylor has in 5 boxes? Taylor's Rocks Number of Boxes Total Number of Rocks 1 7 2 14 3 21 4 28 5 ? aylor organized her rock collection into boxes. She put the same number of rocks into each box. She made this chart to keep track of the total number of rocks she organized. What is the total number of rocks that Taylor has in 5 boxes? Taylor's Rocks Number of Boxes Total Number of Rocks 1 7 2 14 3 21 4 28 5 ? aylor organized her rock collection into boxes. She put the same number of rocks into each box. She made this chart to keep track of the total number of rocks she organized. What is the total number of rocks that Taylor has in 5 boxes? Taylor's Rocks Number of Boxes Total Number of Rocks 1 7 2 14 3 21 4 28 5 ?

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

35

OpenStudy (anonymous):

You just multiply the number of boxes by 7. 1-7 2-14 3-21 4-28 5-35

OpenStudy (anonymous):

@crazychickengirl123 it's 35

OpenStudy (anonymous):

7 times 5 =35

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Welcome medal me please

OpenStudy (anonymous):

@crazychickengirl123 medal me please

OpenStudy (anonymous):

Tn = a + (n-1)d a= 7, n = 5 and d = 7 7 + (5-1)7 T5(i,.e fifth term) = 7 + 4(7)= 7 + 28 =35